Gas is contained in an 8.00-L vessel at a temperature of 20.08 oC and a pressure of 9.00 atm.
(a) Determine the number of moles of gas in the vessel.
(b) How many molecules are in the vessel?
1
Expert's answer
2021-02-15T00:59:12-0500

(a)

"PV=nRT,""n=\dfrac{PV}{RT}=\dfrac{9\ atm\cdot\dfrac{1.013\cdot10^5\ Pa}{1\ atm}\cdot8\ L\cdot\dfrac{1\ m^3}{1000\ L}}{8.314\ \dfrac{J}{K\cdot mol}\cdot293.23\ K}=3\ mol."

(b)

"N=nN_A=3\ mol\cdot6.02\cdot10^{23}\ mol^{-1}=1.81\cdot10^{24}\ molecules."